OpenStudy (anonymous):

HELP! -5/6q-(q-1/2)=1/4(q+1)

OpenStudy (lilg132):

-1/2 = -0.5 1/4 = 0.25 -5/6 = -0.8333333333333333 - * q = -q - * -0.5 = 0.5 so you will have -0.8333333333333333q - q + 0.5 = 0.25q + 0.25 rearrange to get -0.8333333333333333q - q - 0.25q = 0.25 - 0.5 which gives you -2.08333333q = -0.25 -2.08333333q / 2.08333333 = -0.25 / 2.08333333 -q = -0.12 q = 0.12 dont forget to click good answer :)

OpenStudy (lilg132):

or in fraction form q = 3/25